An object has an excess charge of -1.92x10^-7 C.How many excess electrons does it have?

Grade:12

3 Answers

Arun
4 years ago
Dear student
 
Number of excess electrons = -1.92 * 10^-7 / – 1.6 * 10^-19 = 1.2 * 10^12 electrons
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Saurabh Koranglekar
4 years ago
Dear student

Charge = charge of one electron * number of electrons
-1.92*10^(-7)= n (-1.6*10^(-19)

n= 1.2*10^12 electrons

Regards
Khimraj
4 years ago

This object has a charge of 1.92107 C.

Remember that electrons (e) have a charge of 1.6021019 C. So here, there exist: -1.92 * 10-7 / – 1.6 * 10-19 = 1.2*1012